# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### 06 Mar 2026 v4

@@ 5 Classical simulation hardness of shallow-depth boson sampling
As we have previously discussed, since our average-case hardness result considers both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth Boson Sampling as in the original Boson Sampling proposal [aaronson2011computational]. Therefore, in this section, we provide a self-contained analysis of how our average-case hardness result leads to the classical simulation hardness arguments of shallow-depth Boson Sampling. Specifically, we show that if the allowed additive error in Theorem 4 for the hardness is improved to a certain imprecision level, an efficient classical algorithm that can approximately simulate the shallow-depth Boson Sampling is unlikely to exist. This emphasizes that improving the imprecision level of the average-case hardness in Theorem 4 is a crucial step for the classical intractability of shallow-depth Boson Sampling.
Similarly to Refs. [aaronson2011computational, bouland2019complexity], we define an approximate boson sampler as follows.
@@
Given the total variation distance error, the above approximate sampler can have an arbitrarily large additive error for a fixed output probability. Nevertheless, it still has a comparably small additive error for most of the output probabilities due to Markov’s inequality. Accordingly, finding the average-case solution of the output probability of the ideal sampler over randomly chosen collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, up to a certain additive imprecision, is in complexity class $\rm{BPP}^{\rm{NP}}$ by Stockmeyer’s theorem [stockmeyer1985approximation].
Lemma 3 (Average-case approximation [aaronson2011computational]).
@@
We leave the proof of Lemma 3 in Appendix G for a more self-contained analysis. The complexity $\rm{BPP}^{\rm{NP}}$ is known to be inside the finite level of PH, i.e., $\rm{BPP}^{\rm{NP}}\subseteq PH$. Also, by Toda’s theorem [toda1991pp], PH problems can be solved given the ability to solve any #P problem, i.e., $\rm{BPP}^{\rm{NP}}\subseteq\rm{PH}\subseteq\rm{P}^{\rm{\#P}}$. If finding the average-case solution of output probabilities of sampler $\mathcal{S}$ is #P-hard, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Therefore, if an efficient classical algorithm exists that can simulate $\mathcal{S}$, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}}$ which implies the collapse of PH. Consequently, under the assumption of the non-collapse of the PH, there is no efficient classical algorithm capable of simulating $\mathcal{S}$.
@@ 6 Average-case hardness for shallow-depth Gaussian Boson Sampling
Our strategy is to employ the scheme in Ref. [lund2014boson], which used $M_{0}$ product of equally squeezed two-mode squeezed vacuum (TMSV) state as an input state to perform $M_{0}$ mode Boson Sampling task. Specifically, a single TMSV state with squeezing parameter $r$ can be represented as
@@
Note that two $M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture can be embedded in the middle of an $M=2M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture. Accordingly, we define a circuit $C$ in an $M$-mode $\mathcal{B}\mathcal{B}^{*}$ by embedding the given $M_{0}$ mode circuit $C_{0}$ in one side of $\mathcal{B}\mathcal{B}^{*}$, setting gates located right in front of the input ports as balanced beam splitters, and setting the remaining gates as identity gates; we leave in Fig. 3 an illustration of $M=16$ mode circuit $C$ for more clarity. Here, the input $M$ SMSV states with squeezing parameter $r$ combined with the balanced beam splitters at the front becomes $M_{0}$ TMSV states with squeezing parameter $r$. Therefore, our overall setup exactly follows the scheme in Ref. [lund2014boson], such that the first $M_{0}$ mode is the photon counter sector to determine the input configuration of Boson Sampling, and the last $M_{0}$ mode is to simulate the Boson Sampling for the given circuit $C_{0}$.
@@ 7 Hardness analyses in lossy environments
In this section, we generalize our hardness results for $\it{lossy}$ environments, namely, shallow-depth linear optical circuits suffering from photon loss channels after each gate implementation. The reason we consider such a noise channel is that photon loss is indeed a major source of error in optical systems [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian]. Also, photon loss ruins the classical intractability of Boson Sampling, as there exist many efficient classical algorithms that can simulate lossy Boson Sampling within a constant total variation distance [oszmaniec2018classical, garcia2019simulating, qi2020regimes]. Therefore, we mainly deal with the photon loss error here; our goal is to provide evidence for the hardness of the approximate simulation of Boson Sampling in lossy shallow circuits within total variation distance error. For simplicity, we do not consider any photon gain error here, such as thermal radiation noise subjected to the circuits.
To proceed, we start with a brief review of the results presented by Ref. [fujii2016noise], which shows the hardness of simulating noisy quantum circuits. Specifically, one can simulate a noiseless circuit using a larger noisy circuit up to the desired imprecision, by establishing error-detecting code in the noisy circuit and post-selecting null syndrome measurements. Therefore, given the probability to post-select the no-error syndromes, one can approximate the output probability of the noiseless circuit from the output probability of the noisy circuit. Based on this argument, Ref. [bouland2022noise] demonstrates the average-case hardness of approximating output probabilities of noisy quantum circuits, under some plausible assumptions of the noise model. This result gives evidence of the approximate simulation hardness of noisy quantum circuits, within total variation distance error.
@@
where $\mathcal{I}$ is identity, $\mathcal{E}_{i}$ is an CPTP map representing photon loss, and $\rho_{i}$ is a loss rate for each channel satisfying $\rho_{i}\leq\rho$ for a constant $\rho$. The validity of such modeling for photon loss channel is represented in [oszmaniec2018classical, deshpande2022quantum].
@@ 8 Concluding remarks
1. Our result demonstrates the average-case hardness for additive imprecision $2^{-O(N^{\gamma+1}(\log N)^{2})}$. Indeed, there still remains a gap to the desired additive imprecision for the simulation hardness $2^{-(\gamma-1)N\log N-O(N)}$ in Theorem 5. Hence, closing this gap would be an ultimate challenge to the full achievement of classical intractability; more advanced proof techniques are required to reduce this gap. Here, one can take the following approach: finite-size numerical experiments suggest that the output distributions of local random circuits in our circuit architecture (i.e., Definition 2) are close enough to those of global Haar random circuits (see, e.g., Ref. [go2024exploring] and Appendix I). Accordingly, if one can analytically prove that the distance between those output distributions is close enough, we can directly obtain a better imprecision level $2^{-O(N\log N)}$ by results in [bouland2022noise, deshpande2022quantum, bouland2023complexity, bouland2024average], which employed the global Haar random circuits.
@@
Another possible approach for reducing the imprecision gap is to perturb a random circuit matrix in a different way from the Cayley transform (i.e., Definition 8), i.e., as depicted in [bouland2023complexity]. Specifically, instead of perturbing each random gate, one can perturb a submatrix $X$ of our random circuit matrix $U$ with a worst-case matrix $A$ as $X(\theta)=(1-\theta)X+\theta A$ for $\theta\in[0,1]$. Here, a degree of polynomial $|\text{Per}(X(\theta))|^{2}$ is $2N$, which is lower than ours derived by the Cayley transform. Therefore, if one can prove that $X(\theta)$ is distributed similarly to $X$ for small $\theta$, we expect that we can also obtain a better imprecision level by using the same interpolation method. However, the above approach requires one to figure out a global circuit distribution generated by the convolution of local circuit distributions. Although we believe that this problem can be resolved using techniques from random matrix theories, we have not yet developed a complete analysis. Hence, we leave it as an open question.
@@
2. Another important challenge that should be addressed is the extension of the result in Sec. 7, i.e., finding the classical simulation hardness of noisy Gaussian Boson Sampling with photon loss, or Boson Sampling for more general types of noise. As described in [fujii2016noise, bouland2022noise, deshpande2022quantum], employing the threshold theorem would be a viable choice for this goal, which requires efficient error detection codes using only linear optical elements for these setups. However, to the best of our knowledge, such an error detection code does not yet exist. That is, in contrast to the Boson Sampling case, we cannot post-select the so-called ‘no loss’ event in the Gaussian Boson Sampling case because the input photon number is not fixed in this case. This is also for the Boson Sampling with general types of noise because we cannot post-select no error event by using only photon number measurements. Hence, constructing any of these error detection codes would be a crucial step toward the classical hardness of noisy (Gaussian) Boson Sampling, which will contribute to a more noise-tolerant demonstration of quantum advantage with (Gaussian) Boson Sampling. We leave this problem as another open question.
@@
3. One of the existing gaps between theoretical Boson Sampling hardness arguments and Boson Sampling experiments to date is the ratio between mode number $M$ and photon number $N$. Generally, most of the current theoretical Boson Sampling hardness arguments (including our current work) consider only collision-free outcomes and rule out collision outcomes due to the convenience of mathematical analysis [aaronson2011computational, deshpande2022quantum, grier2022complexity, bouland2022noise, bouland2024average]. This calls for the “dilute" regime $M=\Omega(N^{2})$ to ensure the dominance of collision-free outcomes over the probability weight. In contrast, all the experiments to date [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian] operate in the “saturated regime” where the number of modes is comparable to the number of photons. Therefore, to reduce this gap, extending our hardness result of shallow-depth Boson Sampling to the saturated regime $M=O(N^{\gamma})$ for $1\leq\gamma<2$ is crucial, similarly to [bouland2023complexity] for the general Boson Sampling case. To do so, one also needs to consider collision outcomes in the analysis because collision outcomes now dominate the probability weight in this regime. Here, given that our worst-case hardness result in Theorem 3 holds for a linear number of modes $M=\Omega(N)$, this opens the possibility of establishing average-case hardness in the saturated regime by extending the worst-to-average-case reduction from Theorem 3 to include collision outcomes. However, since the collision outcomes do not interchange with each other generally by permutations (in contrast to collision-free outcomes), our proof technique in Sec. 4 using random permutation does not work for general collision outcomes. Therefore, more advanced proof techniques are required to address this challenge. In Appendix H, we present an approach that could offer a useful insight toward extending our average-case hardness result to include collision outcomes. Also, another possible way is to analytically prove that the local random circuit for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ is close enough to those of global Haar random circuits (numerically shown in [go2024exploring] and Appendix I). If this can be done, we can employ the previous argument [bouland2023complexity], which used global Haar random circuits, for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the saturated regime.
@@ Appendix A Previous foundations on average-case hardness
In this appendix, we introduce the existing proof technique used to establish the classical hardness of Boson Sampling, specifically in the context of approximate simulation within a bounded total variation distance error. [aaronson2011computational]. The current state-of-the-art proof technique for the hardness of sampling problems like Boson Sampling essentially builds upon Stockmeyer’s algorithm about approximate counting [stockmeyer1985approximation]. Specifically, given a classical sampler that outputs a sample from a given output distribution, Stockmeyer’s algorithm enables one to multiplicatively estimate a fixed output probability of the sampler, within complexity class $\rm{BPP}^{\rm{NP}}$.
@@
Here, average-case hardness comes into the proof of the classical simulation hardness argument, which means that approximating the ideal output probability of Boson Sampling with high probability over randomly chosen outcomes is #P-hard. More precisely, if the average-case hardness holds up to the imprecision level of the average-case approximation problem, this comes down to the classical simulation hardness of the approximate sampling unless PH collapses, by the complexity-theoretical foundation $\rm{PH}\subseteq\rm{P}^{\#P}$ [toda1991pp].
Moreover, by choosing random circuit instances that have symmetry over the outcomes, one can reduce the average-case instances for the hardness from outcome instances to circuit instances, which is called the hiding property. For the Boson Sampling case, global Haar random unitary (i.e., unitary matrix drawn from Haar measure on U(M), for mode number M) satisfies this condition. In detail, instead of randomly choosing the outcome, we can fix the outcome by applying a random permutation to the global Haar random unitary distribution, which is still Haar distributed due to its translation invariance property. This hiding property plays an important role in the current proofs of the average-case hardness, as it enables one to establish worst-to-average-case reduction in terms of circuit instances. Specifically, as the output probability of Boson Sampling can be written as a low-degree polynomial of input circuit (matrix) values, it allows one to infer the value of a worst-case instance from the output probability of many average-case circuit instances. Hence, the average-case hardness argument for Boson Sampling is typically used in this context, i.e., average-case hardness over random circuit instances, for a fixed outcome [aaronson2011computational, deshpande2022quantum, bouland2022noise].
Accordingly, the crucial step for the classical simulation hardness of approximate sampling is to prove the average-case hardness for the desired imprecision level. While there have been many impressive results and technical developments about the average-case hardness of Boson Sampling [aaronson2011computational, deshpande2022quantum, grier2022complexity, bouland2022noise, bouland2023complexity, bouland2024average], the average-case hardness for the desired imprecision level is not yet fully demonstrated. Still, there exists a gap between the imprecision level of average-case hardness in the strongest existing results and the imprecision level of average-case approximation problem. Hence, closing this imprecision gap remains the ultimate challenge for the complete theoretical demonstration of the classical hardness of approximate Boson Sampling.

### 11 Mar 2026 v5

@@ 5 Classical simulation hardness of shallow-depth boson sampling
As we have previously discussed, since our average-case hardness result considers both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth Boson Sampling as in the original Boson Sampling proposal [1]. Therefore, in this section, we provide a self-contained analysis of how our average-case hardness result leads to the classical simulation hardness arguments of shallow-depth Boson Sampling. Specifically, we show that if the allowed additive error in Theorem 4 for the hardness is improved to a certain imprecision level, an efficient classical algorithm that can approximately simulate the shallow-depth Boson Sampling is unlikely to exist. This emphasizes that improving the imprecision level of the average-case hardness in Theorem 4 is a crucial step for the classical intractability of shallow-depth Boson Sampling.
Similarly to Refs. [1, 41], we define an approximate boson sampler as follows.
@@
Given the total variation distance error, the above approximate sampler can have an arbitrarily large additive error for a fixed output probability. Nevertheless, it still has a comparably small additive error for most of the output probabilities due to Markov’s inequality. Accordingly, finding the average-case solution of the output probability of the ideal sampler over randomly chosen collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, up to a certain additive imprecision, is in complexity class $\rm{BPP}^{\rm{NP}}$ by Stockmeyer’s theorem [48].
Lemma 3 (Average-case approximation [1]).
@@
We leave the proof of Lemma 3 in Appendix G for a more self-contained analysis. The complexity $\rm{BPP}^{\rm{NP}}$ is known to be inside the finite level of PH, i.e., $\rm{BPP}^{\rm{NP}}\subseteq PH$. Also, by Toda’s theorem [49], PH problems can be solved given the ability to solve any #P problem, i.e., $\rm{BPP}^{\rm{NP}}\subseteq\rm{PH}\subseteq\rm{P}^{\rm{\#P}}$. If finding the average-case solution of output probabilities of sampler $\mathcal{S}$ is #P-hard, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Therefore, if an efficient classical algorithm exists that can simulate $\mathcal{S}$, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}}$ which implies the collapse of PH. Consequently, under the assumption of the non-collapse of the PH, there is no efficient classical algorithm capable of simulating $\mathcal{S}$.
@@ 6 Average-case hardness for shallow-depth Gaussian Boson Sampling
Our strategy is to employ the scheme in Ref. [50], which used $M_{0}$ product of equally squeezed two-mode squeezed vacuum (TMSV) state as an input state to perform $M_{0}$ mode Boson Sampling task. Specifically, a single TMSV state with squeezing parameter $r$ can be represented as
@@
Note that two $M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture can be embedded in the middle of an $M=2M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture. Accordingly, we define a circuit $C$ in an $M$-mode $\mathcal{B}\mathcal{B}^{*}$ by embedding the given $M_{0}$ mode circuit $C_{0}$ in one side of $\mathcal{B}\mathcal{B}^{*}$, setting gates located right in front of the input ports as balanced beam splitters, and setting the remaining gates as identity gates; we leave in Fig. 3 an illustration of $M=16$ mode circuit $C$ for more clarity. Here, the input $M$ SMSV states with squeezing parameter $r$ combined with the balanced beam splitters at the front becomes $M_{0}$ TMSV states with squeezing parameter $r$. Therefore, our overall setup exactly follows the scheme in Ref. [50], such that the first $M_{0}$ mode is the photon counter sector to determine the input configuration of Boson Sampling, and the last $M_{0}$ mode is to simulate the Boson Sampling for the given circuit $C_{0}$.
@@ 7 Hardness analyses in lossy environments
In this section, we generalize our hardness results for $\it{lossy}$ environments, namely, shallow-depth linear optical circuits suffering from photon loss channels after each gate implementation. The reason we consider such a noise channel is that photon loss is indeed a major source of error in optical systems [4, 5, 6, 7]. Also, photon loss ruins the classical intractability of Boson Sampling, as there exist many efficient classical algorithms that can simulate lossy Boson Sampling within a constant total variation distance [15, 16, 17]. Therefore, we mainly deal with the photon loss error here; our goal is to provide evidence for the hardness of the approximate simulation of Boson Sampling in lossy shallow circuits within total variation distance error. For simplicity, we do not consider any photon gain error here, such as thermal radiation noise subjected to the circuits.
To proceed, we start with a brief review of the results presented by Ref. [51], which shows the hardness of simulating noisy quantum circuits. Specifically, one can simulate a noiseless circuit using a larger noisy circuit up to the desired imprecision, by establishing error-detecting code in the noisy circuit and post-selecting null syndrome measurements. Therefore, given the probability to post-select the no-error syndromes, one can approximate the output probability of the noiseless circuit from the output probability of the noisy circuit. Based on this argument, Ref. [28] demonstrates the average-case hardness of approximating output probabilities of noisy quantum circuits, under some plausible assumptions of the noise model. This result gives evidence of the approximate simulation hardness of noisy quantum circuits, within total variation distance error.
@@
where $\mathcal{I}$ is identity, $\mathcal{E}_{i}$ is an CPTP map representing photon loss, and $\rho_{i}$ is a loss rate for each channel satisfying $\rho_{i}\leq\rho$ for a constant $\rho$. The validity of such modeling for photon loss channel is represented in [15, 3].
@@ 8 Concluding remarks
1. Our result demonstrates the average-case hardness for additive imprecision $2^{-O(N^{\gamma+1}(\log N)^{2})}$. Indeed, there still remains a gap to the desired additive imprecision for the simulation hardness $2^{-(\gamma-1)N\log N-O(N)}$ in Theorem 5. Hence, closing this gap would be an ultimate challenge to the full achievement of classical intractability; more advanced proof techniques are required to reduce this gap. Here, one can take the following approach: finite-size numerical experiments suggest that the output distributions of local random circuits in our circuit architecture (i.e., Definition 2) are close enough to those of global Haar random circuits (see, e.g., Ref. [47] and Appendix I). Accordingly, if one can analytically prove that the distance between those output distributions is close enough, we can directly obtain a better imprecision level $2^{-O(N\log N)}$ by results in [28, 3, 29, 30], which employed the global Haar random circuits.
@@
Another possible approach for reducing the imprecision gap is to perturb a random circuit matrix in a different way from the Cayley transform (i.e., Definition 8), i.e., as depicted in [29]. Specifically, instead of perturbing each random gate, one can perturb a submatrix $X$ of our random circuit matrix $U$ with a worst-case matrix $A$ as $X(\theta)=(1-\theta)X+\theta A$ for $\theta\in[0,1]$. Here, a degree of polynomial $|\text{Per}(X(\theta))|^{2}$ is $2N$, which is lower than ours derived by the Cayley transform. Therefore, if one can prove that $X(\theta)$ is distributed similarly to $X$ for small $\theta$, we expect that we can also obtain a better imprecision level by using the same interpolation method. However, the above approach requires one to figure out a global circuit distribution generated by the convolution of local circuit distributions. Although we believe that this problem can be resolved using techniques from random matrix theories, we have not yet developed a complete analysis. Hence, we leave it as an open question.
@@
2. Another important challenge that should be addressed is the extension of the result in Sec. 7, i.e., finding the classical simulation hardness of noisy Gaussian Boson Sampling with photon loss, or Boson Sampling for more general types of noise. As described in [51, 28, 3], employing the threshold theorem would be a viable choice for this goal, which requires efficient error detection codes using only linear optical elements for these setups. However, to the best of our knowledge, such an error detection code does not yet exist. That is, in contrast to the Boson Sampling case, we cannot post-select the so-called ‘no loss’ event in the Gaussian Boson Sampling case because the input photon number is not fixed in this case. This is also for the Boson Sampling with general types of noise because we cannot post-select no error event by using only photon number measurements. Hence, constructing any of these error detection codes would be a crucial step toward the classical hardness of noisy (Gaussian) Boson Sampling, which will contribute to a more noise-tolerant demonstration of quantum advantage with (Gaussian) Boson Sampling. We leave this problem as another open question.
@@
3. One of the existing gaps between theoretical Boson Sampling hardness arguments and Boson Sampling experiments to date is the ratio between mode number $M$ and photon number $N$. Generally, most of the current theoretical Boson Sampling hardness arguments (including our current work) consider only collision-free outcomes and rule out collision outcomes due to the convenience of mathematical analysis [1, 3, 27, 28, 30]. This calls for the “dilute" regime $M=\Omega(N^{2})$ to ensure the dominance of collision-free outcomes over the probability weight. In contrast, all the experiments to date [4, 5, 6, 7] operate in the “saturated regime” where the number of modes is comparable to the number of photons. Therefore, to reduce this gap, extending our hardness result of shallow-depth Boson Sampling to the saturated regime $M=O(N^{\gamma})$ for $1\leq\gamma<2$ is crucial, similarly to [29] for the general Boson Sampling case. To do so, one also needs to consider collision outcomes in the analysis because collision outcomes now dominate the probability weight in this regime. Here, given that our worst-case hardness result in Theorem 3 holds for a linear number of modes $M=\Omega(N)$, this opens the possibility of establishing average-case hardness in the saturated regime by extending the worst-to-average-case reduction from Theorem 3 to include collision outcomes. However, since the collision outcomes do not interchange with each other generally by permutations (in contrast to collision-free outcomes), our proof technique in Sec. 4 using random permutation does not work for general collision outcomes. Therefore, more advanced proof techniques are required to address this challenge. In Appendix H, we present an approach that could offer a useful insight toward extending our average-case hardness result to include collision outcomes. Also, another possible way is to analytically prove that the local random circuit for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ is close enough to those of global Haar random circuits (numerically shown in [47] and Appendix I). If this can be done, we can employ the previous argument [29], which used global Haar random circuits, for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the saturated regime.
@@ Appendix A Previous foundations on average-case hardness
In this appendix, we introduce the existing proof technique used to establish the classical hardness of Boson Sampling, specifically in the context of approximate simulation within a bounded total variation distance error. [1]. The current state-of-the-art proof technique for the hardness of sampling problems like Boson Sampling essentially builds upon Stockmeyer’s algorithm about approximate counting [48]. Specifically, given a classical sampler that outputs a sample from a given output distribution, Stockmeyer’s algorithm enables one to multiplicatively estimate a fixed output probability of the sampler, within complexity class $\rm{BPP}^{\rm{NP}}$.
@@
Here, average-case hardness comes into the proof of the classical simulation hardness argument, which means that approximating the ideal output probability of Boson Sampling with high probability over randomly chosen outcomes is #P-hard. More precisely, if the average-case hardness holds up to the imprecision level of the average-case approximation problem, this comes down to the classical simulation hardness of the approximate sampling unless PH collapses, by the complexity-theoretical foundation $\rm{PH}\subseteq\rm{P}^{\#P}$ [49].
Moreover, by choosing random circuit instances that have symmetry over the outcomes, one can reduce the average-case instances for the hardness from outcome instances to circuit instances, which is called the hiding property. For the Boson Sampling case, global Haar random unitary (i.e., unitary matrix drawn from Haar measure on U(M), for mode number M) satisfies this condition. In detail, instead of randomly choosing the outcome, we can fix the outcome by applying a random permutation to the global Haar random unitary distribution, which is still Haar distributed due to its translation invariance property. This hiding property plays an important role in the current proofs of the average-case hardness, as it enables one to establish worst-to-average-case reduction in terms of circuit instances. Specifically, as the output probability of Boson Sampling can be written as a low-degree polynomial of input circuit (matrix) values, it allows one to infer the value of a worst-case instance from the output probability of many average-case circuit instances. Hence, the average-case hardness argument for Boson Sampling is typically used in this context, i.e., average-case hardness over random circuit instances, for a fixed outcome [1, 3, 28].
Accordingly, the crucial step for the classical simulation hardness of approximate sampling is to prove the average-case hardness for the desired imprecision level. While there have been many impressive results and technical developments about the average-case hardness of Boson Sampling [1, 3, 27, 28, 29, 30], the average-case hardness for the desired imprecision level is not yet fully demonstrated. Still, there exists a gap between the imprecision level of average-case hardness in the strongest existing results and the imprecision level of average-case approximation problem. Hence, closing this imprecision gap remains the ultimate challenge for the complete theoretical demonstration of the classical hardness of approximate Boson Sampling.